Session: Nov 1998 (session 4)
Wavelength: 6 cm
To give feedback on this experiment, click
Medicina (Franco Mantovani at Mon Nov 23 04:13:44 MET 1998) Success
Noto (C.Trigilio at Mon Nov 23 10:21:16 MET 1998) Success
Torun (K.Borkowski at Mon Nov 23 10:51:03 MET 1998) Success
One of the BBCs (No 07) was unstable (too low signal) till 20:55 UTC.
Medicina ( at Mon Nov 23 11:52:05 MET 1998) Success
Effelsberg ( at Mon Nov 23 11:52:21 MET 1998) Success
Westerbork (Tony Foley at Tue Nov 24 16:31:24 MET 1998) Minor failures
Basically looks OK, but one of the 8 cryogenic frontends had warmed up.
Jodrell (Alastair Gunn at Wed Nov 25 07:48:38 MET 1998) Success
No known problems.
Cambridge (Alastair Gunn at Wed Nov 25 08:19:19 MET 1998) Minor failures
Tape failed to start on first pass. Appeared to be a problem
with the terminal buffer. Reset FS and data good from 20:04 UT.
Onsala (A. Polatidis at Wed Nov 25 16:21:17 MET 1998) Success
Shanghai (Huang Xinyong at Fri Nov 27 06:45:08 MET 1998) Success
The first scan is fail, because of the form was error in equip.ctl.
The FS restart before scond scan. It is OK.
Shanghai (Huang Xinyong at Fri Nov 27 06:53:39 MET 1998) Success
The first scan is fail, because of the form error in equip.ctl. FS
restart before second scan. It is OK.
Calibration (Jean\-Francois Desmurs at Fri Dec 11 16:19:39 MET 1998) Minor failures
(Bad values for JB and no value for CM).
The ANTAB formatted file for this experiment can be found on VLBEER
in the nov98 directory under the name:
es023.antab. Read the comments in the file to get further